# Post-hoc Stochastic Concept Bottleneck Models

### v2

@@ Abstract
Concept Bottleneck Models (CBMs) are interpretable models that predict the target variable through high-level human-understandable concepts, allowing users to intervene on mispredicted concepts to adjust the final output. While recent work has shown that modeling dependencies between concepts can improve CBM performance, especially under interventions, such approaches typically require retraining the entire model, which may be infeasible when access to the original data or compute is limited. In this paper, we introduce Post-hoc Stochastic Concept Bottleneck Models (PSCBMs), a lightweight method that augments any pre-trained CBM with a multivariate normal distribution over concepts by adding only a small covariance-prediction module, without retraining the backbone model. We propose two training strategies and show on real-world data that PSCBMs consistently match or improve both concept and target accuracy over standard CBMs at test time. Furthermore, we show that due to the modeling of concept dependencies, PSCBMs perform much better than CBMs under interventions, while remaining far more efficient than retraining a similar stochastic model from scratch.
@@ 1 Introduction
The adoption of machine learning models in high-stakes domains is often limited by their lack of interpretability due to their black-box nature (lipton2016mythos; doshi-velez2017towards). Concept Bottleneck Models (CBMs), first introduced by koh2020concept, address this by inserting a layer of neurons aligned with human-understandable concepts before the target prediction. A CBM is composed of a concept encoder that predicts the concepts, and a prediction head that takes these concepts as inputs; the final prediction is thus explained through the intermediate concept values. While the original formulation assumes independence among concepts, accounting for correlations has been shown to improve the model’s performance (havasi2022addressing; singhi2024improvinginterventionefficacyconcept; Vandenhirtz2024stochastic). However, existing approaches that capture concept dependencies typically require training the entire model with dedicated objectives. In contrast, we demonstrate that such dependencies can be incorporated post-hoc into a pre-trained CBM. A more detailed discussion of related work is provided in Appendix˜A.
Building on this idea, we introduce Post-Hoc Stochastic Concept Bottleneck Models (PSCBMs), which extend pre-trained CBMs with a lightweight module that predicts concept covariance. Inspired by Stochastic Concept Bottleneck Models (SCBMs) (Vandenhirtz2024stochastic), we model concept dependencies with a multivariate normal distribution. This approach not only improves predictive performance, but also enables more efficient concept interventions—the process by which a user modifies predicted concepts to directly influence downstream predictions.
This work contributes to model interpretability through human-understandable concepts in three ways: (i) We introduce PSCBM, a lightweight post-hoc module that incorporates concept correlations into existing CBMs without requiring full retraining, thereby reducing data and compute requirements. (ii) We propose a simple intervention-based training procedure that further improves intervention efficiency in (P)SCBM-like models. (iii) We show on real-world data that PSCBM improves both predictive accuracy and intervention effectiveness while remaining substantially more efficient than full model retraining.
@@ 2 Method
Figure 1: Comparison of CBM (left), SCBM (right), and our proposed PSCBM (right, red block). All methods input $\bm{x}$} to an encoder to produce the feature vector $\bm{z}$} from which concepts are obtained and fed to the target predictor. CBM directly predicts concept values $\hat{{\bm{c}}}$, while in SCBM, the predictor outputs an expected value $\bm{\mu}$ and a covariance matrix $\bm{\Sigma}$ that define a multivariate normal distribution to sample from. PSCBM incorporates the $\bm{\Sigma}$ predictor (red box) to a pre-trained CBM.
@@
In this paper we propose PSCBMs [^1] The code is available at https://github.com/wiktorhof/PSCBM., a post-hoc extension of CBMs that enables to model dependencies between concepts without retraining the model from scratch. A CBM koh2020concept consists of three main components: a feature encoder, a concept predictor, and a target predictor. The feature encoder $h$ (e.g., a CNN or ResNet) extracts features ${\bm{z}}=h({\bm{x}})$ from an input $\bm{x}$}. Then the concept predictor $g$ outputs concept probabilities ${\bm{p}}=g({\bm{z}})$. When focusing on binary concepts for downstream prediction (i.e. hard CBMs (havasi2022addressing)), rather than their probabilities (soft CBMs), the next step is sampling $M$ concept values from a Bernoulli distrbution $\hat{{\bm{c}}}^{(1)},\dots\hat{{\bm{c}}}^{(M)}\sim\text{Bernoulli}({\bm{p}})$. Then the target predictor $f$ maps the sampled concepts to the final prediction $\hat{y}=\frac{1}{M}\sum_{m=1}^{M}f(\hat{{\bm{c}}}^{(m)})$. SCBMs (Vandenhirtz2024stochastic) extend this framework by explicitly capturing correlations between concepts: the concept predictor is extended to feature a mean predictor $g_{\mu}$ and a covariance predictor $g_{\Sigma}$, which define a multivariate normal distribution that models concept dependencies. Concepts are sampled as $\hat{{\bm{c}}}^{(m)}=\sigma(\eta^{(m)})$ for $m=1,\dots,M$, where $\eta^{(m)}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ and $\sigma$ is the sigmoid function. Our proposed PSCBMs directly build upon this idea: given a pre-trained CBM, we reuse its concept predictor $g$ as $g_{\mu}$ and add a lightweight covariance predictor $g_{\Sigma}$, training only the latter while keeping the rest of the model frozen. In this way, any existing CBM can be turned into a stochastic, dependency-aware model post-hoc, as illustrated in Figure˜1.
@@
We propose two training paradigms for PSCBM, which can be extended to SCBM: (i) The loss function is applied to the model’s predictions without concept interventions. (ii) We encourage more responsive predictors to concept interventions. In every training iteration, for every sample in the training dataset, we randomly select a set $\mathcal{S}$ of concepts to intervene on using some strategy $\tau$ and calculate the loss after the intervention. The cardinality of $\mathcal{S}$ is fixed throughout training. This is done $N$ times for each data point and the average of the loss for all these interventions is calculated. This method is similar to RandInt introduced by espinosa2022concept but it differs in two ways: (i) Unlike RandInt, which decides independently for each concept whether to replace it with its true value, our method enforces that a fixed number of concepts is intervened on in every case. (ii) At each training iteration, multiple random interventions per sample are made and the loss is averaged. This reduces gradient variance and yields more stable training compared to applying a single intervention. A detailed pseudocode for the calculation of this loss can be found in Appendix˜C.
@@ 3 Results
We compare the performance of PSCBM to SCBM and regular CBM. PSCBM indicates training without concept interventions, while PSCBMi includes interventions during training. SCBM proposes two ways to learn the covariance matrix, amortized per instance, $\bm{\Sigma}({\bm{x}})$, or learnt globally, $\bm{\Sigma}$. For brevity, we present the results on the global covariance, and include their amortized counterparts in Appendix˜D. All implementation details are described in Appendix˜E. For test-time interventions, concepts are selected based on their associated prediction uncertainty, i.e. the concept with predicted probability closest to 0.5 is chosen. Unlike CBM where the intervened probabilities are set to 0/1, in SCBMs and PSCBMs concept logits are set to thus of $\varepsilon$ or $1-\varepsilon$, respectively, where $\varepsilon$ is small for stability. We evaluate our methods on the Caltech-UCSD Birds-200-2011 dataset (wah2011caltech), composed of photographs of birds of 200 different classes. We use the variant introduced in koh2020concept for CBMs with 112 binary per-class concepts and, as evaluation metrics, we focus on concept and target accuracy.
@@
To evaluate intervention performance we rely on two approaches: visual inspection of the intervention curves, and the area under the intervention curve (AUC) (singhi2024improvinginterventionefficacyconcept), which summarizes each curve into a single value for easier comparison. The last two columns of Table˜1 report the AUC scores. SCBM achieves the highest concept AUC, but both PSCBM variants outperform it on target AUC, with PSCBMi being the strongest overall. This highlights the benefit of training with interventions. Both PSCBM models also outperform the regular CBM on both metrics, showing that adding a covariance matrix post-hoc improves intervention performance. For a more detailed comparison, we examine the intervention trajectories under the uncertainty policy in Figure˜2. We additionally include the corresponding results when selecting concepts with a random policy in Appendix˜F. Both PSCBM variants consistently outperform CBM, with the intervention-trained variant performing best. However, for a few interventions, PSCBM without interventions during training lags behind SCBM, indicating that joint training of the covariance is more effective. PSCBM with interventions surpasses SCBM on target accuracy after about 20 interventions, though it does not reach SCBM’s performance on concept accuracy. We note that SCBM could also be trained with interventions, which might further boost its results. It is clear from these that adding a covariance on a trained CBM can significantly improve its performance.
@@ Appendix A Related Work
Concept Bottleneck Models (CBM) have originally been proposed by koh2020concept. The idea is to transform a regular neural network into an interpretable model by converting one of its layers into a concept bottleneck, where each neuron corresponds to a high-level concept that can be interpreted by a human. Such a model can be framed as a composite function $f({\bm{x}})=g(h({\bm{x}}))$, where $h(\cdot)$ is a concept predictor that calculates the concepts $\hat{{\bm{c}}}$, and $g(\cdot)$ is the target predictor that predicts the final label from the concepts. Since all information that goes from the input to the output passes through the concept bottleneck, the network’s prediction should be well explained by the concepts. koh2020concept use what has been called soft concept encoding: the target predictor uses real-valued concept logits or probabilities produced by the encoder. However, it has been shown that this method is prone to information leakage, where undesired information can pass through the bottleneck [margeloiu2021do, mahinpei2021promises, makonnen2025measuring]. As a remedy, havasi2022addressing propose hard concepts: the concept encoder predicts concept probabilities, which parametrize Bernoulli distributions, from which binary inputs for the target predictor are sampled. Additionally, they introduce in their work a side channel which allows the target predictor to use information that cannot be expressed in terms of the high-level concepts. This modification can increase target accuracy if the concept set is incomplete. An alternative approach to modeling the concept bottleneck has been proposed by espinosa2022concept who introduce Concept Embedding Models (CEM), where each concept is modeled not by a single neuron but instead by a pair of learnable vector embeddings corresponding to the presence or absence of a concept in the input. This allows passing information not captured by the concepts through the bottleneck without the need for a side channel. Several methods have been proposed to transform regular models into Concept Bottleneck Models. yuksekgonul2023post propose a method to convert one layer of a neural network trained without concepts into a concept bottleneck by using only a small concept-annotated subset. oikarinen2023label demonstrate that concepts can also be generated with the help of a Vision-Language Model such as CLIP [radford2021learning], reducing the human effort associated with annotating data. This work focuses on vision tasks but can be further extended to other domains as applications of CBMs have been explored as well in the context of text [laguna2025interpretable] or tabular data [zarlengatabcbm].
@@
A particular advantage of CBMs is the possibility of affecting downstream predictions by modifying incorrectly predicted concepts. This process is called a concept intervention. koh2020concept only consider interventions on randomly selected concepts. On the other hand, chauhan2023interactive, shin2023closer propose and evaluate more efficient concept selection policies, such as choosing the concept with the highest associated prediction uncertainty, that is, with probability closest to 0.5. This policy allows reducing target prediction error by half after 12 concept interventions on average, while with random concept selection an average of 43 interventions is necessary to achieve it. In espinosa2022concept, a regularization technique called RandInt is introduced, whereby randomly selected concepts are set to their true value during training. This is then shown to increase the model’s responsiveness to test-time concept interventions. As an alternative for an explicit concept selection policy, espinosa2024learning, singhi2024improvinginterventionefficacyconcept parameterize it by a neural network which can be trained jointly with the model or in a post-hoc manner. A further extension of the interventions framework is done by laguna2024beyond, introducing a method for intervening on a black-box model without altering its architecture. They also define a measure of intervenability and show that fine-tuning a model for this quantity can improve the performance of interventions.
@@
Since correlated concepts may require multiple interventions, several methods model dependencies. havasi2022addressing use an autoregressive structure, by which the value of every concept depends on the previous concepts. It increases the model’s accuracy, but has the disadvantage that concepts have to be evaluated sequentially, which increases evaluation time. singhi2024improvinginterventionefficacyconcept augment the CBM with a neural network that is trained to update all concepts in response to an intervention made to one of them. This can be combined with the trainable concept selection policy mentioned above. These elements can be either trained jointly with the full model or added to an existing CBM post-hoc. However, unlike our method, it does not provide an explicit representation of concept correlations. Finally, Vandenhirtz2024stochastic introduce Stochastic Concept Bottleneck Models (SCBM), in which concept correlations are captured by a multivariate normal distribution. It requires predicting not only the expected value of the concepts, but also a covariance matrix. A special advantage of this model is the explicit representation of concept dependencies, which is used in an intervention strategy based on the distribution’s confidence region.

### v4

@@ Abstract
Concept Bottleneck Models (CBMs) are interpretable models that predict the target variable through high-level human-understandable concepts, allowing users to intervene on mispredicted concepts to adjust the final output. While recent work has shown that modeling dependencies between concepts can improve CBM performance, especially under interventions, such approaches typically require retraining the entire model, which may be infeasible when access to the original data or compute is limited. In this paper, we introduce Post-hoc Stochastic Concept Bottleneck Models (PSCBMs), a lightweight method that augments any pre-trained CBM with a multivariate normal distribution over concepts by adding only a small covariance-prediction module, without modifying the backbone model. We propose two training strategies and show on real-world data that PSCBMs consistently match or improve both concept and target accuracy over standard CBMs at test time. Furthermore, we show that due to the modeling of concept dependencies, PSCBMs perform much better than CBMs under interventions, while remaining far more efficient than retraining a similar stochastic model from scratch.
@@ 1 Introduction
The adoption of machine learning models in high-stakes domains is often limited by their lack of interpretability due to their black-box nature (Lipton, 2016; Doshi-Velez and Kim, 2017). Concept Bottleneck Models (CBMs), first introduced by Koh et al. (2020), address this by inserting a layer of neurons aligned with human-understandable concepts before the target prediction. A CBM is composed of a concept encoder that predicts the concepts, and a prediction head that takes these concepts as inputs; the final prediction is thus explained through the intermediate concept values. While the original formulation assumes independence among concepts, accounting for correlations has been shown to improve the model’s performance (Havasi et al., 2022; Singhi et al., 2024; Vandenhirtz et al., 2024). However, existing approaches that capture concept dependencies typically require training the entire model with dedicated objectives. In contrast, we demonstrate that such dependencies can be incorporated post-hoc into a pre-trained CBM. A more detailed discussion of related work is provided in Appendix˜A.
Building on this idea, we introduce Post-Hoc Stochastic Concept Bottleneck Models (PSCBMs) [^1] The code is available at https://github.com/wiktorhof/PSCBM., which extend pre-trained CBMs with a lightweight module that predicts concept covariance. Inspired by Stochastic Concept Bottleneck Models (SCBMs) (Vandenhirtz et al., 2024), we model concept dependencies with a multivariate normal distribution. This approach not only improves predictive performance, but also enables more efficient concept interventions—the process by which a user modifies predicted concepts to directly influence downstream predictions.
This work contributes to model interpretability through human-understandable concepts in three ways: (i) We introduce PSCBM, a lightweight post-hoc module that incorporates concept correlations into existing CBMs without requiring full retraining, thereby reducing compute requirements. (ii) We propose a simple intervention-based training procedure that further improves intervention efficiency in (P)SCBM-like models. (iii) We show on real-world data that PSCBM improves both predictive accuracy and intervention effectiveness while remaining substantially more efficient than full model retraining.
@@ 2 Method
Figure 1: Comparison of a hard CBM (top), SCBM (bottom), and our proposed PSCBM (bottom, red block). All methods input $\bm{x}$} to an encoder to produce the feature vector $\bm{z}$} from which concepts are obtained and fed to the target predictor. CBM directly predicts concept values $\hat{{\bm{c}}}$, while in SCBM, the predictor outputs an expected value $\bm{\mu}$ and a covariance matrix $\bm{\Sigma}$ that define a multivariate normal distribution to sample from. PSCBM incorporates the $\bm{\Sigma}$ predictor (red box) to a frozen pre-trained CBM.
@@
In this paper we propose PSCBMs, a post-hoc extension of CBMs that enables to model dependencies between concepts without retraining the model from scratch. A CBM Koh et al. (2020) consists of three main components: a feature encoder, a concept predictor, and a target predictor. The feature encoder $h$ (e.g., a CNN or ResNet) extracts features ${\bm{z}}=h({\bm{x}})$ from an input $\bm{x}$}. Then the concept predictor $g$ outputs concept probabilities ${\bm{p}}=g({\bm{z}})$. When focusing on binary concepts for downstream prediction (i.e. hard CBMs (Havasi et al., 2022)), rather than their probabilities (soft CBMs), the next step is sampling $M$ concept values from a Bernoulli distrbution $\hat{{\bm{c}}}^{(1)},\dots\hat{{\bm{c}}}^{(M)}\sim\text{Bernoulli}({\bm{p}})$. Then the target predictor $f$ maps the sampled concepts to the final prediction $\hat{y}=\frac{1}{M}\sum_{m=1}^{M}f(\hat{{\bm{c}}}^{(m)})$. SCBMs (Vandenhirtz et al., 2024) extend this framework by explicitly capturing correlations between concepts: the concept predictor is extended to feature a mean predictor $g_{\mu}$ and a covariance predictor $g_{\Sigma}$, which define a multivariate normal distribution that models concept dependencies. Concepts are sampled as $\hat{{\bm{c}}}^{(m)}=\sigma(\eta^{(m)})$ for $m=1,\dots,M$, where $\eta^{(m)}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ and $\sigma$ is the sigmoid function. Our proposed PSCBMs directly build upon this idea: given a pre-trained CBM, we reuse its concept predictor $g$ as $g_{\mu}$ and add a lightweight covariance predictor $g_{\Sigma}$, training only the latter while keeping the rest of the model frozen. In this way, any existing CBM can be turned into a stochastic, dependency-aware model post-hoc, as illustrated in Figure˜1.
@@
We propose two training paradigms for PSCBM, which can be extended to SCBM: (i) The loss function is applied to the model’s predictions without concept interventions. (ii) We encourage more responsive predictors to concept interventions. In every training iteration, for every sample in the training dataset, we randomly select a set $\mathcal{S}$ of concepts to intervene on using some strategy $\tau$ and calculate the loss after the intervention. The cardinality of $\mathcal{S}$ is fixed throughout training. This is done $N$ times for each data point and the average of the loss for all these interventions is calculated. This method is similar to RandInt introduced by Espinosa Zarlenga et al. (2022) but it differs in two ways: (i) Unlike RandInt, which decides independently for each concept whether to replace it with its true value, our method enforces that a fixed number of concepts is intervened on in every case. (ii) At each training iteration, multiple random interventions per sample are made and the loss is averaged. This reduces gradient variance and yields more stable training compared to applying a single intervention. A detailed pseudocode for the calculation of this loss can be found in Appendix˜C.
@@ 3 Results
We compare the performance of PSCBM to SCBM and regular CBM. PSCBM indicates training without concept interventions, while PSCBMi includes interventions during training. SCBM proposes two ways to learn the covariance matrix, amortized per instance, $\bm{\Sigma}({\bm{x}})$, or learnt globally, $\bm{\Sigma}$. For brevity, we present the results on the global covariance, and include their amortized counterparts in Appendix˜D. All implementation details are described in Appendix˜E. For test-time interventions, concepts are selected based on their associated prediction uncertainty, i.e., the concept with predicted probability closest to 0.5 is chosen. Unlike CBM where the intervened probabilities are set to 0/1, SCBMs and PSCBMs set intervened concept logits to those corresponding to $\epsilon$ or $1-\epsilon$, respectively, where $\varepsilon$ is small for stability. We evaluate our methods on the Caltech-UCSD Birds-200-2011 dataset (Wah et al., 2011), composed of photographs of birds of 200 different classes. We use the variant introduced in Koh et al. (2020) for CBMs with 112 binary per-class concepts and, as evaluation metrics, we focus on concept and target accuracy.
@@
To evaluate intervention performance we rely on two approaches: visual inspection of the intervention curves, and the area under the intervention curve (AUC) (Singhi et al., 2024), which summarizes each curve into a single value for easier comparison. The last two columns of Table˜1 report the AUC scores. SCBM achieves the highest concept AUC, but both PSCBM variants outperform it on target AUC, with PSCBMi being the strongest overall. This highlights the benefit of training with interventions. Both PSCBM models also outperform the regular CBM on both metrics, showing that adding a covariance matrix post-hoc improves intervention performance. For a more detailed comparison, we examine the intervention trajectories under the uncertainty policy in Figure˜2. We additionally include the corresponding results when selecting concepts with a random policy in Appendix˜F. Both PSCBM variants consistently outperform CBM, with the intervention-trained variant performing best. However, for a few interventions, PSCBM without interventions during training lags behind SCBM, indicating that joint training of the covariance is more effective. PSCBM with interventions surpasses SCBM on target accuracy after about 20 interventions, though it does not reach SCBM’s performance on concept accuracy. We note that SCBM could also be trained with interventions, which might further boost its results. It is clear from these that adding a covariance on a trained CBM can significantly improve its performance.
@@ Appendix A Related Work
Concept Bottleneck Models (CBM) have originally been proposed by Koh et al. (2020). The idea is to transform a regular neural network into an interpretable model by converting one of its layers into a concept bottleneck, where each neuron corresponds to a high-level concept that can be interpreted by a human. Such a model can be framed as a composite function $f({\bm{x}})=g(h({\bm{x}}))$, where $h(\cdot)$ is a concept predictor that calculates the concepts $\hat{{\bm{c}}}$, and $g(\cdot)$ is the target predictor that predicts the final label from the concepts. Since all information that goes from the input to the output passes through the concept bottleneck, the network’s prediction should be well explained by the concepts. Koh et al. (2020) use what has been called soft concept encoding: the target predictor uses real-valued concept logits or probabilities produced by the encoder. However, it has been shown that this method is prone to information leakage, where undesired information can pass through the bottleneck (Margeloiu et al., 2021; Mahinpei et al., 2021; Makonnen et al., 2025). As a remedy, Havasi et al. (2022) propose hard concepts: the concept encoder predicts concept probabilities, which parametrize Bernoulli distributions, from which binary inputs for the target predictor are sampled. Additionally, they introduce in their work a side channel which allows the target predictor to use information that cannot be expressed in terms of the high-level concepts. This modification can increase target accuracy if the concept set is incomplete. An alternative approach to modeling the concept bottleneck has been proposed by Espinosa Zarlenga et al. (2022) who introduce Concept Embedding Models (CEM), where each concept is modeled not by a single neuron but instead by a pair of learnable vector embeddings corresponding to the presence or absence of a concept in the input. This allows passing information not captured by the concepts through the bottleneck without the need for a side channel. Several methods have been proposed to transform regular models into Concept Bottleneck Models. Yuksekgonul et al. (2023) propose a method to convert one layer of a neural network trained without concepts into a concept bottleneck by using only a small concept-annotated subset. Oikarinen et al. (2023) demonstrate that concepts can also be generated with the help of a Vision-Language Model such as CLIP (Radford et al., 2021), reducing the human effort associated with annotating data. This work focuses on vision tasks but can be further extended to other domains as applications of CBMs have been explored as well in the context of text (Laguna et al., 2025) or tabular data (Zarlenga et al., ).
@@
A particular advantage of CBMs is the possibility of affecting downstream predictions by modifying incorrectly predicted concepts. This process is called a concept intervention. Koh et al. (2020) only consider interventions on randomly selected concepts. On the other hand, Chauhan et al. (2023); Shin et al. (2023) propose and evaluate more efficient concept selection policies, such as choosing the concept with the highest associated prediction uncertainty, that is, with probability closest to 0.5. This policy allows reducing target prediction error by half after 12 concept interventions on average, while with random concept selection an average of 43 interventions is necessary to achieve it. In Espinosa Zarlenga et al. (2022), a regularization technique called RandInt is introduced, whereby randomly selected concepts are set to their true value during training. This is then shown to increase the model’s responsiveness to test-time concept interventions. As an alternative for an explicit concept selection policy, Espinosa Zarlenga et al. (2024); Singhi et al. (2024) parameterize it by a neural network which can be trained jointly with the model or in a post-hoc manner. A further extension of the interventions framework is done by Laguna et al. (2024), introducing a method for intervening on a black-box model without altering its architecture. They also define a measure of intervenability and show that fine-tuning a model for this quantity can improve the performance of interventions.
@@
Since correlated concepts may require multiple interventions, several methods model dependencies. Havasi et al. (2022) use an autoregressive structure, by which the value of every concept depends on the previous concepts. It increases the model’s accuracy, but has the disadvantage that concepts have to be evaluated sequentially, which increases evaluation time. Singhi et al. (2024) augment the CBM with a neural network that is trained to update all concepts in response to an intervention made to one of them. This can be combined with the trainable concept selection policy mentioned above. These elements can be either trained jointly with the full model or added to an existing CBM post-hoc. However, unlike our method, it does not provide an explicit representation of concept correlations. Finally, Vandenhirtz et al. (2024) introduce Stochastic Concept Bottleneck Models (SCBM), in which concept correlations are captured by a multivariate normal distribution. It requires predicting not only the expected value of the concepts, but also a covariance matrix. A special advantage of this model is the explicit representation of concept dependencies, which is used in an intervention strategy based on the distribution’s confidence region.